In a rectangular coordinate system a charge of 25x10 C is placed at the origin of coordinates, and a charge of-25x10-9 C is placed at the point-6m, y:0. What is the electric field at (a) x-3 m, y-0,
Answer #1

a) At x = 3m, y=0. 32 9 o 3m 25x10-9C 13m I - 25 x 10-96 Ep = K (25x10-9, ·x2 tp. = 9x25x2 i > Ep, = 50 NCE Hence Enet = 50